Abraham is 8 years older than Belen. The sum of Belen's age and Celeste's age is three years more than twice Abraham's age. If the sum of their three ages is 138, how old are Abraham, Belen & Celeste?
Answer by macston(5194)

.
.
A=Abraham; B=Belen; C=Celeste
A=B-8yrs
B+C=2A+3yrs Substitute for A
B+C=2(B-8yrs)+3yrs
B+C=2B-16yrs+3yrs Subtract B from each side.
C=B-13yrs
A+B+C=138 yrs Substitute for A and C
(B-8yrs)+B+(B-13 yrs)=138yrs
3B-21yrs=138yrs Add 21 yrs to each side.
3B=159 yrs Divide each side by 3.
B=53 yrs ANSWER 1: Belen is 53 years old.
A=B-8yrs=53yrs-8yrs=45 yrs ANSWER 2: Abraham is 45 years old.
C=B-13yrs=53yrs-13yrs=40 yrs ANSWER 3: Celeste is 40 years old.
CHECK:
A+B+C=138 years
45yrs+53yrs+40yrs=138yrs
138yrs=138yrs